OpenStudy (anonymous):

Given the function f(x) = 12x + 9, evaluate f(12) .a.)96 b.)11.3 c.)21 d.)153

OpenStudy (anonymous):

where ever there is x, substitute 12 in it.. so it would be 12(12)+9.. so that would be....?

OpenStudy (lncognlto):

=12(12) + 9 =144 + 9 = 153

OpenStudy (anonymous):

thank you both very much :)